# Kairos: Towards Adaptive and Generalizable Time Series Foundation Models

### v1

@@ 3.5 Training Corpus
For pre-training, we curated Predictability-Stratified Time Series (PreSTS) corpus, a corpus of over 300 billion time points. PreSTS comprises real-world time series spanning multiple domains and is augmented by a synthetic dataset we constructed to ensure comprehensive coverage. Additionally, we have excluded all test sets found within the GIFT-Eval benchmark (Aksu et al., 2024) to maintain evaluation integrity. While a prevailing consensus regarding TSFMs highlights the necessity of a diverse training dataset, research has also shown that the presence of anomalies in the training set can severely degrade a model’s predictive capabilities (Cheng et al., 2024). Therefore, the real-world datasets in PreSTS are stratified into five tiers based on their degree of predictability. Datasets exhibiting higher predictability are assigned a greater sampling probability during training, a strategy that enables the model to undergo a more robust, high-quality training regimen without sacrificing the comprehensiveness of the data. Comprehensive descriptions of our training datasets along with synthetic data generation algorithm, are provided in Appendix E.1.
@@ 4 Experiment
In this section, we conduct a comprehensive evaluation to assess the performance of Kairos and the effect of proposed components. Specifically, we investigate the following three key research questions. RQ1: Does Kairos achieve better generalization and zero-shot prediction than previous methods? (Section 4.2) RQ2: How effectively does MoS-DP handle time series heterogeneity by adapting its patch size according to local information densities? (Section 4.3 and 4.4.1) RQ3: Does IARoPE effectively generate customized position embedding tailored to the unique temporal structure of each time series instance? (Section 4.3 and 4.4.2)
@@ 4.1.1 Evaluation Datasets and Metric
To comprehensively evaluate the performance of our proposed method, we conducted zero-shot assessments on two well-known public benchmarks: GIFT-Eval (Aksu et al., 2024) and Time-Series-Library (TSLib) (Wang et al., 2024b). We train Kairos in three sizes: mini (10M), small (23M) and base (50M). The training details are provided in Appendix C.1.
Benchmarks. GIFT-Eval benchmark consists of 97 prediction tasks, with 55 tasks dedicated to short-term predictions, 21 for medium-term predictions, and 21 for long-term predictions, providing a comprehensive means of assessing the model’s predictive capabilities. For the TSLib benchmark, we followed the common practice in Time-MoE (Shi et al., 2024) and selected the ETT datasets (Zhou et al., 2021a) and Weather (Zhou et al., 2021b). Additionally, to test the model’s adaptability to diverse sampling frequencies, we incorporated the Saugeen datasets (Godahewa et al., 2021), which include data with both daily and weekly sampling frequencies. These were referred to as Saugeen (D) and Saugeen (W), respectively. The detailed description of the evaluation datasets, their splitting configurations, and the selection of the evaluation length are provided in Appendix E.2 and E.3.
Metric. In our evaluation of GIFT-Eval, consistent with prior research (Auer et al., 2025; Liu et al., 2025), we employ the official Mean Absolute Scaled Error (MASE) and Continuous Ranked Probability Score (CRPS) metrics to assess point and probabilistic forecasting performance, respectively. These metrics are first normalized by the Seasonal Naïve baseline for each individual task. The final score is then computed as the geometric mean of these normalized values. For the TSLib benchmark, we followed previous works (Liu et al., 2023; 2024c; Nie et al., 2023) by adopting Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate time series forecasting performance. The metric calculations are detailed in Appendix C.2.1.
@@ 4.1.2 Compared Methods
We compare Kairos with several state-of-the-art models, including TSFMs such as Sundial (Liu et al., 2025), Toto (Cohen et al., 2025), YingLong (Wang et al., 2025), VisionTS (Chen et al., 2025), Time-MoE (Shi et al., 2024), ChronosBolt (Ansari et al., 2024), TTM (Ekambaram et al., 2024), Moirai (Woo et al., 2024), Timer-XL (Liu et al., 2024c), TimesFM-2.0 (Das et al., 2024), and Chronos (Ansari et al., 2024), as well as advanced full-shot deep learning models, including PatchTST (Nie et al., 2023), DLinear (Zeng et al., 2022), iTransformer (Liu et al., 2023), Pathformer (Chen et al., 2024), and TimesNet (Wu et al., 2022). .
@@ 4.2 Zero-shot Evaluation
Finding 1: Kairos demonstrates remarkable zero-shot forecasting capabilities across multiple benchmarks. On the GIFT-Eval benchmark, as presented in Table 1, $\textsc{Kairos}_{\text{base}}$ (50M) achieves the best MASE and the second-best CRPS when compared to state-of-the-art task-specific deep learning methods and other TSFMs. Notably, $\textsc{Kairos}_{\text{small}}$ (23M) surpasses both Toto and Sundial in MASE, despite having a parameter count that is 6.6 and 5.6 times smaller, respectively. On the TSLib benchmark, as shown in Figure 3, our lightweight $\textsc{Kairos}_{\text{mini}}$ (10M) surpasses the performance of both recent advanced TSFMs and the majority of full-shot deep learning models. Compared with other TSFMs, Kairos achieves a significant improvement in predictive performance.
Table 1: Performance evaluation on the GIFT-Eval. We evaluated model performance using the normalized MASE and CRPS metrics delineated in Section 4.1.1, where lower values indicate higher forecasting accuracy. The baseline models fall into three categories: statistical methods, deep learning (DL) models, and TSFMs. Following the official classification of GIFT-Eval, TSFMs are further subdivided into TestData Leakage and Zero-Shot. TestData Leakage signifies that the model’s training set included test data, whereas Zero-Shot indicates that the model was trained without any data from the test set or its corresponding training split. Baseline results are officially reported by GIFT-Eval.
Method
Seasonal
Naïve
DLinear
PTST. TTM Chronos
Chronos
Bolt
TimesFM Moirai VisionTS Ying. Toto Sundial
$\textsc{Kairos}_{s}$
$\textsc{Kairos}_{b}$
#Params - - - 5M 709M 205M 500M 311M 112M 300M 151M 128M 23M 50M
MASE 1.000 1.061 0.849 1.020 0.870 0.808 0.758 0.875 0.863 0.798 0.750 0.750 0.748 0.742
CRPS 1.000 0.846 0.587 0.873 0.574 0.574 0.550 0.599 0.755 0.548 0.517 0.559 0.554 0.548
Figure 3: Zero-shot forecasting performance on TSLib. Results are averaged across prediction lengths {96, 192, 336, 720}. The subscripts $l$, $b$, $s$, and $a$ represent model sizes of large, base, small, and advanced, respectively. The complete experimental results are presented in Appendix F.
@@ 4.3 Ablation Study
Table 2: Ablation study comparing MoS-DP and RoPE variants. We evaluated ablation settings using the normalized MASE from Section 4.1.1, assessing performance on individual prediction horizons and in aggregate across all tasks.
MoS-DP RoPE Variant Short Medium Long AVG
IARoPE 0.719 0.759 0.789 0.742
IARoPE 0.725 0.779 0.813 0.755
RoPE 0.726 0.783 0.828 0.759
RoPE 0.731 0.784 0.829 0.763
Finding 2: MoS-DP and IARoPE collaboratively enable Kairos to adapt to heterogeneous time series, achieving powerful predictive performance. To investigate the effectiveness of our proposed MoS-DP and IARoPE,we conducted a comprehensive ablation study, with the results summarized in Table 2. For a fair comparison, Kairos without MoS-DP uses a fixed patch size of 64, which was determined to be optimal via hyperparameter search. For IARoPE, we compare against the original RoPE (Su et al., 2024), which employs a fixed, predefined $\theta$. The integration of MoS-DP and IARoPE into our method leads to the best performance on the 97 tasks of the GIFT-Eval benchmark, which validates the superiority of our designs. Furthermore, the improvement of Kairos gets more significant compared to Kairos without MoS-DP or IARoPE as horizon gets longer, exhibiting better long-term forecasting ability.
@@ 4.4 Model Analysis
4.4.1 Patch Distribution Analysis
Finding 3: Kairos demonstrates adaptive capability to varying information densities, enabling it to model complex temporal dynamics with appropriate granularity. For each coarsest patch, Kairos dynamically fuses information by selecting from a set of granularities (patch sizes). To quantify this dynamic routing, we calculate a weighted average patch size by multiplying the weight assigned to each granularity by its corresponding patch size. Consequently, a smaller value for this metric indicates the model’s preference for finer-grained processing of a given coarsest patch, while a larger value suggests a preference for coarser granularity. In Figure 4, this weighted average patch size is visualized by the background color, with darker shades corresponding to smaller values. As the figure illustrates, Kairos strategically employs smaller patch sizes for sequences with abrupt changes and larger patch sizes for stable sequences.
Figure 4: Patch size preferences in GIFT-Eval test datasets. Darker shades indicate a smaller weighted average patch size, signifying the model’s preference for finer-grained processing in that region.
4.4.2 IARoPE Analysis
Figure 5: Causal analysis of adaptive modulation by IARoPE. This experiment validates the criticality of matching positional encodings to the unique characteristics of each time series instance by disrupting or removing this adaptation. We test this by manipulating the RoPE frequencies $\theta$ under several conditions: IARoPE (standard), Intra-Dataset Shuffle ($\theta$ modulations permuted between different instances within the same dataset), Inter-Dataset Shuffle ($\theta$ modulations from instances of other datasets), and Fixed RoPE (no modulation).
Finding 4: Modulating RoPE frequencies $\theta$ instance-wise can indeed better model time series temporal relationships. To verify that the performance gains of IARoPE are indeed caused by its instance-specific RoPE frequencies $\theta$ modulation, we designed a series of control experiments that impair or ablate this mechanism by shuffling the modulation parameters (details in Figure 5; Appendix E.4). The results confirm IARoPE’s superior forecasting accuracy, demonstrating the advantage of tailoring PE to individual instance frequency characteristics. The Intra-Dataset Shuffle ($\theta$ modulations permuted between different instances within the same dataset) performed second best, suggesting that while instance-specific modulation is optimal, leveraging dataset-level frequency similarities still offers utility. Conversely, both Inter-Dataset Shuffle ($\theta$ modulations from instances of other datasets) and Fixed RoPE (no modulation) yielded poorer performance, highlighting that mismatched or static $\theta$ parameters are suboptimal as they fail to capture the specific temporal nature of each series.
@@ 5 Conclusions
We introduce Kairos, a time series foundation model that advances the design of modules specifically adapted to the unique characteristics of time series data. Kairos achieves this through key innovations, including Mixture-of-Size Dynamic Patching, which considers varying information densities for more precise and efficient modeling, and Instance-Adaptive RoPE, which provides more accurate positional information tailored to time series nuances. Our comprehensive experiments reveal superior performance achieved by Kairos, highlighting the importance of designing foundational models that are inherently aligned with the structural nuances of time series data.
Appendix A Analysis of Multi-Patch Prediction
@@
Appendix B Analysis of Inference Speed
@@
TTM-Advanced* 0.012
Timer-XL 0.021
ChronosBolt-Base 0.062
Kairos (Ours) 0.065
Moirai-Large 0.077
Time-MoE-Large 0.156
TimesFM 0.210
@@
Appendix C Implementation Details
C.1 Training Details
C.1.1 Model Configurations
We train Kairos in three sizes: mini (10M), small (23M) and base (50M) with the detailed model configurations are in Table 4. The base model are trained for 300,000 steps with batch sizes of 512. We employ the AdamW optimizer, a linear decay learning rate adjustment strategy for model optimization. The learning rate for parameters related to IARoPE is set to 1e-5, while the learning rate for others is set to 1e-3. Training is conducted on 4 × NVIDIA A100 GPUs using TF32 precision, which takes only 15 hours for base size.

### v2

@@ 3.5 Training Corpus
For pre-training, we curated the Predictability-Stratified Time Series (PreSTS) corpus, comprising over 300 billion time points. PreSTS integrates large-scale real-world time series from diverse domains and a complementary synthetic dataset to ensure broad coverage. All test sets from the GIFT-Eval (Aksu et al., 2024) and TSLib (Wang et al., 2024b) benchmarks are explicitly excluded to prevent data leakage. While dataset diversity is widely regarded as essential for TSFM training, recent studies indicate that low-predictability or anomalous sequences can substantially impair forecasting (Cheng et al., 2024). To address this, we stratify real-world datasets into five tiers based on predictability and assign higher sampling probabilities to more predictable data during training. This strategy promotes stable and high-quality pre-training without compromising data diversity. Further details are provided in Appendix E.1.
@@ 4 Experiment
In this section, we assess the performance of Kairos and its components by addressing three research questions. RQ1: Does Kairos achieve superior zero-shot generalization capabilities while maintaining parameter efficiency compared to previous methods? (Section 4.2) RQ2: Does the Mixture-of-Size Encoder effectively handle time series heterogeneity by adapting its granularity based on local information density? (Section 4.3 and 4.4.1) RQ3: Does DRoPE effectively generate customized position embeddings tailored to the unique temporal structure of each time series instance? (Section 4.3 and 4.4.2)
Table 1: Performance evaluation on the GIFT-Eval. We evaluate using normalized MASE and CRPS (Section 4.1.1). Baselines encompass statistical, Deep Learning (DL), and TSFM methods, with TSFMs categorized into TestData Leakage and Zero-Shot following official protocols. TestData Leakage signifies that the model’s training set included test data, whereas Zero-Shot indicates that the model was trained without any data from the test set or its corresponding training split. Baseline results are officially reported by GIFT-Eval.
Method Seasonal Naïve DLinear PTST. TTM Chronos Chronos Bolt TimesFM Moirai VisionTS Ying. Toto Sundial $\textsc{Kairos}_{s}$ (ours) $\textsc{Kairos}_{b}$ (ours)
#Params - - - 5M 709M 205M 500M 311M 112M 300M 151M 128M 23M 53M
MASE $\downarrow$ 1.000 1.061 0.849 1.020 0.870 0.808 0.758 0.875 0.863 0.798 0.750 0.750 0.748 0.738
CRPS $\downarrow$ 1.000 0.846 0.587 0.873 0.574 0.574 0.550 0.599 0.755 0.548 0.517 0.559 0.554 0.548
Figure 3: Zero-shot performance on TSLib averaged over prediction lengths {96, 192, 336, 720}. The subscripts $l$, $b$, $s$, $m$, and $a$ represent model sizes of large, base, small, mini and advanced, respectively. The complete experimental results are presented in Appendix F.
@@ 4.1.1 Evaluation Datasets and Metric
To comprehensively evaluate our proposed method, we conducted zero-shot assessments on two well-known public benchmarks: GIFT-Eval (Aksu et al., 2024) and Time-Series-Library (TSLib) (Wang et al., 2024b). We train Kairos in three sizes: mini (10M), small (23M) and base (53M). The training details are provided in Appendix D.1.
Benchmarks. The GIFT-Eval benchmark comprises 97 tasks spanning short- (55), medium- (21), and long-term (21) horizons, facilitating a comprehensive assessment. For TSLib, we selected the ETT and Weather (Zhou et al., 2021) datasets, consistent with Time-MoE (Shi et al., 2024). To test adaptability to diverse frequencies, we also included the Saugeen datasets (Godahewa et al., 2021), denoted as Saugeen (D) and Saugeen (W) for daily and weekly intervals, respectively. The detailed description of the evaluation datasets, splitting configurations, and evaluation length selection are provided in Appendix E.2 and E.3.
Metric. In our evaluation of GIFT-Eval, consistent with prior research (Auer et al., 2025; Liu et al., 2025), we employ the official Mean Absolute Scaled Error (MASE) and Continuous Ranked Probability Score (CRPS) metrics to assess point and probabilistic forecasting performance, respectively. These metrics are first normalized by the Seasonal Naïve baseline for each individual task. The final score is then computed as the geometric mean of these normalized values. For the TSLib benchmark, we followed previous works (Liu et al., 2023; 2024c; Nie et al., 2023) by adopting Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate time series forecasting performance. The metric calculations are detailed in Appendix D.2.1.
@@ 4.1.2 Compared Methods
We compare Kairos with several state-of-the-art models, including TSFMs such as Sundial (Liu et al., 2025), Toto (Cohen et al., 2025), YingLong (Wang et al., 2025), VisionTS (Chen et al., 2025), Time-MoE (Shi et al., 2024), ChronosBolt (Ansari et al., 2024), TTM (Ekambaram et al., 2024), Moirai (Woo et al., 2024), Timer-XL (Liu et al., 2024c), TimesFM-2.0 (Das et al., 2024), Chronos (Ansari et al., 2024), and advanced full-shot deep learning models, including PatchTST (Nie et al., 2023), DLinear (Zeng et al., 2022), iTransformer (Liu et al., 2023), Pathformer (Chen et al., 2024), and TimesNet (Wu et al., 2022).
@@ 4.2 Zero-shot Evaluation
Finding 1: Kairos achieves superior zero-shot forecasting performance with significantly higher parameter efficiency. On the GIFT-Eval benchmark, as presented in Table 1, $\textsc{Kairos}_{\text{base}}$ (53M) achieves the best MASE and the second-best CRPS among state-of-the-art task-specific deep learning methods and other TSFMs. Notably, $\textsc{Kairos}_{\text{small}}$ (23M) surpasses both Toto and Sundial in MASE, despite having a parameter count that is 6.6 and 5.6 times smaller, respectively. On the TSLib benchmark, as shown in Figure 3, our lightweight $\textsc{Kairos}_{\text{mini}}$ (10M) surpasses the performance of both recent advanced TSFMs and the majority of full-shot deep learning models. These results confirm that Kairos delivers a significant improvement in predictive capability without relying on massive scale.
@@ 4.3 Ablation Study
Finding 2: The synergistic integration of Mixture-of-Size Encoding, DRoPE, and the Multi-Patch Decoding is critical for adapting to heterogeneous time series and achieving superior forecasting performance. Table 2 details the component-wise contributions. (i) Encoder Design: Kairos outperforms variants using fixed patch sizes or excluding null experts, confirming that dynamic, sparse multi-granularity modeling is critical for capturing diverse temporal patterns. (ii) Positional Embedding: DRoPE proves superior to standard RoPE (Su et al., 2024). Our ablation further isolates the gains from Instance-Level Frequency Modulation and Granularity-Aware Position Calibration, verifying the benefit of adapting to both spectral heterogeneity and physical time distortion caused by dynamic patching. (iii) Decoding Strategy: The Multi-Patch Decoding shows clear superiority over standard single-patch autoregressive decoding. The integration of both components yielded the best performance across 97 GIFT-Eval tasks, which validates the superiority of our designs. Furthermore, the performance gap widens as the forecast horizon increases, demonstrating Kairos’s superior long-term forecasting capability.
Table 2: Ablation study comparing encoder design, positional embedding and decoder strategy. Models were evaluated using the normalized MASE (detailed in Section 4.1.1) across individual prediction horizons and as an aggregate across all tasks.
Model Variants Short Medium Long AVG
Impact of Encoder Design
   w/ Fixed Patch Size ($P=32$) 0.724 0.802 0.820 0.761
   w/ Mixture-of-Size (w/o Null Experts) 0.720 0.770 0.800 0.748
Impact of Positional Embedding
   w/ Standard RoPE (Fixed $\theta$) 0.729 0.807 0.835 0.767
   w/ Only Instance-Level Frequency Modulation 0.719 0.767 0.797 0.746
   w/ Only Granularity-Aware Position Calibration 0.727 0.796 0.807 0.758
Impact of Decoder Strategy
   w/ Single-Patch Autoregressive 0.705 0.794 0.848 0.753
[…]
@@ 4.4 Model Analysis
4.4.1 Mixture-of-Size Tokenization Analysis
Figure 4: Adaptive tokenization visualization. Tokens are demarcated by dotted lines and varying colors, where deeper red curves indicate finer granularity. Kairos adapts granularity to information density, applying finer processing to volatile regions.
Finding 3: Kairos demonstrates adaptive capability to varying information densities, enabling it to model complex temporal dynamics with appropriate granularity. Instead of enforcing a uniform resolution, Kairos dynamically routes segments to optimal granularity based on local information. As visualized in Figure 4, Kairos strategically allocates fine-grained tokens to regions with abrupt changes or high volatility, while employing coarse-grained tokens for stable trends. This adaptive allocation mechanism reduces computational redundancy on low-information segments, allowing the model to achieve superior performance with a compact parameter budget compared to rigid architectures. See Appendix C for a detailed inference speed analysis.
4.4.2 DRoPE Analysis
Table 3: Causal analysis of adaptive modulation by DRoPE. We evaluate the impact of disrupting or removing the instance-wise $\theta$ adjustment. Degradation indicates relative performance drop compared to DRoPE.
Method DRoPE Intra-Dataset Standard RoPE Inter-Dataset
(Ours) Shuffle (Fixed $\theta$) Shuffle
MASE $\downarrow$ 0.738 0.751 0.767 0.947
Degradation -1.76% -3.93% -28.32%
Finding 4: Modulating RoPE frequencies $\theta$ instance-wise can indeed better model time series temporal relationships. To verify that DRoPE’s performance gains stem from its instance-specific $\theta_{\text{inst}}$ modulation, we designed control experiments that disrupt this mechanism (Table 3; Appendix E.4): (i) Intra-Dataset Shuffle: randomly permute $\theta_{\text{inst}}$ across instances within the same dataset; (ii) Inter-Dataset Shuffle: assign $\theta_{\text{inst}}$ from instances of different datasets; (iii) Fixed RoPE: use standard RoPE without any modulation. The results reveal a clear performance gradient (DRoPE $\succ$ Intra-Shuffle $\succ$ Fixed RoPE $\succ$ Inter-Shuffle), indicating that DRoPE learns instance-specific representations to modulate $\theta$. Intra-Dataset Shuffle outperforms Fixed RoPE, as samples within the same dataset share similar spectral characteristics, allowing shuffled $\theta_{\text{inst}}$ to remain partially compatible. However, Intra-Dataset Shuffle still incurs a 1.76% degradation compared to DRoPE, suggesting that instance-level adaptation contributes to the performance gains beyond dataset-level statistics. Inter-Dataset Shuffle yields the worst results (28.32% degradation), as $\theta_{\text{inst}}$ from different datasets cannot capture the target series’ temporal dynamics. These results collectively validate the effectiveness of instance-level $\theta_{\text{inst}}$ adaptation in DRoPE.
@@ 5 Conclusions
We introduce Kairos, a parameter-efficient foundation model explicitly tailored for time series heterogeneity. Its architecture features three key innovations: a Mixture-of-Size Encoder for adaptive granularity, a Heterogeneity-Aware Transformer with Dynamic RoPE for instance-specific temporal calibration, and a Multi-Patch Decoder for efficient parallel forecasting. Extensive experiments demonstrate that Kairos achieves superior performance, validating the necessity of aligning architectural inductive biases with the structural nuances of time series data. One limitation of our current approach is that Kairos relies on channel-independent modeling (Nie et al., 2023) to support multivariate time series and does not explicitly capture inter-variable dependencies. Future work aims to model these interactions and extend Kairos to tasks beyond forecasting.
Impact Statement
Our work on Kairos is foundational research focused on advancing parameter-efficient and adaptive time series modeling. We anticipate several positive societal impacts: (i) promoting sustainable AI by significantly reducing the computational resources and energy required for large-scale deployment; (ii) enhancing decision-making in dynamic environments (e.g., energy, retail) through improved modeling of heterogeneous temporal patterns; and (iii) providing a high-quality data curation methodology via the PreSTS corpus to improve model performance.
While we do not foresee immediate negative societal consequences, we recognize that advanced predictive technologies carry inherent risks of misuse. Consequently, rigorous ethical oversight is therefore essential to guarantee that such predictive capabilities serve the public interest.
Appendix A Etymology of Kairos
Kairos was chosen for its mythological significance as the Greek god of the “right or critical moment”, reflecting our model’s ability to select the ideal set of granularities and positional encoding for time series characterized by varying information densities and frequencies. Furthermore, this continues a thematic tradition seen in other foundational models we benchmark against, such as Chronos (Ansari et al., 2024) (the personification of time) and Moirai (Woo et al., 2024) (the personifications of destiny).
Appendix B Analysis of Multi-Patch Prediction
Unlike task-specific, non-autoregressive models such as TFT (Lim et al., 2021) and Informer (Zhou et al., 2021), which predict all future time points simultaneously, their architecture confines them to forecasting a fixed length identical to the training sequence, precluding variable-length predictions. Conversely, TSFMs typically employ an iterative process: they predict a segment of the sequence, append it to the historical data, and then forecast the subsequent segment, thereby enabling autoregressive predictions of arbitrary length. To mitigate the cumulative errors inherent in this autoregressive methodology, Kairos introduces forecast tokens designed to predict multiple patches within each autoregressive step.
@@ Appendix B Analysis of Multi-Patch Prediction
Appendix C Analysis of Inference Speed
@@
TTM-Advanced* 0.009
Timer-XL 0.014
ChronosBolt-Base 0.055
Kairos-Base (Ours) 0.061
Moirai-Large 0.070
Sundial-Base 0.099
Time-MoE-Large 0.148
TimesFM-2.0 0.202
@@
Toto-Base 13.717
Appendix D Implementation Details
D.1 Training Details
D.1.1 Model Configurations
We train Kairos in three sizes: mini (10M), small (23M) and base (53M) with the detailed model configurations are in Table 5. The base model are trained for 300,000 steps with batch sizes of 512. We employ the AdamW optimizer, a linear decay learning rate adjustment strategy for model optimization. The learning rate for parameters related to DRoPE is set to 1e-5, while the learning rate for others is set to 1e-3. Training is conducted on 4 × NVIDIA A100 GPUs using TF32 precision, which takes only 15 hours for base size.